# Back from the dead: AT2019aalc as a candidate repeating TDE in an AGN

### v2

@@ Abstract
Aims. We aim to classify the transient and explain the mechanism responsible for its second optical flare.
@@
Results. The observations revealed a uniquely bright UV counterpart and multiple X-ray flares during the second optical flaring episode of the transient. The second flare, similarly to the first one, is also accompanied by IR dust echo emission. A long-term radio flare is found with an inverted spectrum. Optical spectroscopic observations reveal the presence of Bowen Fluorescence lines and strong high-ionization coronal lines indicating an extreme level of ionization in the system.
Conclusions. The results suggest that the transient can be classified as a Bowen Fluorescence Flare (BFF), a relatively new sub-class of flaring active galactic nuclei (AGN). AT2019aalc can be also classified as an extreme coronal line emitter (ECLE). We found that, in addition to AT2019aalc, another BFF AT2021loi is spatially coincident with a high-energy neutrino event. The multi-wavelength properties of these transients suggest a possible connection between ECLEs, BFFs and TDEs in AGN.
@@
Tidal disruption events (TDEs) are rare transient events, which occur when a star approaches a supermassive black hole (SMBH) and the tidal forces of the latter rip the star apart (Rees 1988). About half of the star’s material is accreted around the black hole, generating a luminous outburst across the electromagnetic spectrum. TDE optical light curves are characterized by quick and colorless evolving but also fast (on monthly timescales) decaying emission. Several TDEs were detected in X-rays with a very soft spectrum typically explained by late-time accretion disk formation (Saxton et al. 2020). In addition, a handful of TDEs produced detectable radio emission, explained by delayed111Note that the SMBH accretion rate does not peak on the same timescale as the mass fall-back rate. non-relativistic outflows (Horesh et al. 2021; Cendes et al. 2023), or only in instances by newly launched relativistic on-axis jets (Berger et al. 2012; Brown et al. 2015; Cenko et al. 2012; De Colle & Lu 2020; Yao et al. 2023a). Spectroscopically, TDEs are characterized by a strong blue continuum and broad emission lines ( km s-1) e.g., H and H. In a few cases, Bowen Fluorescence (BF: Bowen 1934, 1935) lines e.g., N iii at appear (e.g., Blagorodnova et al. 2019; Charalampopoulos et al. 2022a).
@@
TDEs generally result in one optical flare, however, in a few cases a re-brightening episode was detected, even years after the initial flare. These might be explained in the following way. Under special circumstances, a star on a grazing orbit might be partially disrupted and only a fraction of the stellar mass will be deposited onto the SMBH. The secondary disruption of a surviving core after it reaches the pericenter again results in re-brightening. Only a few partial TDEs or candidates are known to date. These are AT2020vdq (Somalwar et al. 2023), ASASSN 14ko (Payne et al. 2021), eRASSt J045650.3-203750 (Liu et al. 2023), RX J133157.6-324319.7 (Hampel et al. 2022), AT2018fyk (Wevers et al. 2019, 2023), AT2019avd (Chen et al. 2022), AT2022dbl (Lin et al. 2024) and AT2021aeuk (Sun et al. 2025).
@@
Based on the observed properties of TDE host galaxies, French et al. (2020) estimated a TDE rate of per year per galaxy which is well borne out by recent data (Yao et al. 2023b). Importantly, a similar fraction of TDEs should take place in AGN hosts (e.g., Chan et al. 2019; Ryu et al. 2024). These transients are challenging to distinguish from regular AGN variability. Only a few candidate TDEs taken place in AGN (in the following: TDE-AGN) were studied earlier (e.g., PS16dtm - a TDE in a narrow-line Seyfert-1 (NLSy1) galaxy (Blanchard et al. 2017) or the multiple soft X-ray flares in IC 3599 (Campana et al. 2015) and GSN 069 (Shu et al. 2018)). In contrast with TDEs seen in earlier quiescent galaxies, these transients are characterized by slowly decaying optical emission and bumps tend to appear after the peak.
@@
The family tree of AGN-related transients became much broader in the past years (e.g., Frederick et al. 2021). In addition to TDEs, flaring NLSy1s, changing-look AGN (CLAGN) and different types of SNe were classified as typical nuclear transients. Recently, Trakhtenbrot et al. (2019) identified a new class of spectroscopically unique flares from accreting SMBHs. In , the transient events AT2017bgt, F01004-2237 and OGLE17aaj have been identified as the first members of the class based on their unique optical spectroscopic features. Later, the peculiar X-ray transient AT2019avd (Trakhtenbrot et al. 2020) was found with remarkable common properties, however, its classification as an AT2017bgt-like transient is unclear. A more promising case is AT2021loi (Makrygianni et al. 2023), a UV-bright transient event with very similar multi-wavelength properties to the instances studied by Trakhtenbrot et al. (2019). These events show unobscured AGN-like optical spectra with significant and persistent BF lines, extremely strong UV emission and slowly, on yearly scales, decaying optical emission. These AGN are characterized by significantly and longer-term intensified accretion. Hereafter, we will refer to this new class of SMBH flares as Bowen Fluorescence Flares (BFFs), following Makrygianni et al. (2023). Using the Zwicky Transient Facility (ZTF, Masci et al. 2019) Public Survey data, Dgany et al. (2023) found only one BFF candidate (AT2021seu) among spectroscopically classified transients, suggesting that these flares are rare. Two other candidates, namely AT2022fpx, (Koljonen et al. 2024) and AT2020afhd (Arcavi et al. 2024) were found later.
@@
In recent years, two nuclear transients were found to be spatially coincident with neutrino events detected by the IceCube Neutrino Observatory based on ZTF follow-up campaigns of IceCube neutrino alerts (Stein et al. 2023). The TDE AT2019dsg is likely associated with the neutrino, IC191001A (Stein et al. 2021) and a further promising association, between the candidate TDE AT2019fdr and IC200530A (Reusch et al. 2022) points to TDEs as a potential new class of sources of the extragalactic neutrinos. Both sources were identified as having luminous mid-infrared emission, and a search for similar flares (van Velzen et al. 2024) led to the identification of the transient event AT2019aalc (a.k.a. ZTF19aaejtoy) with the neutrino alert IC191119A. In addition, two candidate obscured TDEs (Jiang et al. 2023) and the extremely energetic candidate TDE AT2021lwx (Yuan et al. 2024) were associated with high-energy neutrinos. A recently claimed association between a neutrino flare detected by the IceCube Observatory and an X-ray flare from the TDE ATLAS17jrp suggests that the X-ray emission of these sources play a crucial rule in producing neutrinos (Li et al. 2024).
@@
In this paper, we focus on the nuclear transient event and neutrino source candidate AT2019aalc especially on its significant optical re-brightening that has started years after its initial flare (Veres et al. 2023).
@@
The paper is organized as follows. We explain the discovery of the transient event and give an insight about its host galaxy in Sect. 2, summarize our observing campaign and the data reduction in Sect. 3, present the results of the observations in Sect. 4 and discuss them in Sect. 5 and finally give a summary and conclude in Sect. 6. Throughout the paper, we adopt a flat CDM cosmological model with parameters , , and . In this model (Wright 2006), mas angular size corresponds to pc projected linear size at the source redshift (Ahn et al. 2012). This redshift corresponds to a luminosity distance of Mpc. All magnitudes are given in AB system (Oke 1974).
@@ 2 The transient event: AT2019aalc
An analysis of the infrared properties of AT2019dsg and AT2019fdr revealed a strong reverberation signal i.e. the infrared dust echo, known as reprocessed emission of optical/UV photons in the infrared by hot dust around the SMBH, at a distance of pc (Jiang et al. 2016; Dou et al. 2016; Lu et al. 2016; van Velzen et al. 2021b; Jiang et al. 2021; van Velzen et al. 2024). In addition to the exceptionally high infrared luminosities of these transients, the neutrino detection times appear to be temporally consistent with the infrared luminosity peaks (Stein et al. 2021; Reusch et al. 2022). This discovery motivated a systematic search for neutrino emission using an extended sample of black hole flares (van Velzen et al. 2024). This archival search revealed the nuclear transient AT2019aalc with a powerful optical flare and dust echo emission in spatial coincidence with the high-energy neutrino event, IC191119A (IceCube Collaboration 2019) detected by the IceCube Observatory.
The host galaxy of AT2019aalc (2MASX J15241665+0451192) is a barred spiral (SBa) galaxy hosting an active nucleus (Oh et al. 2013). Based on its Sloan Digital Sky Survey (SDSS: Abazajian et al. 2009) spectrum it was further classified as a broad-line Seyfert-1 galaxy (Trump et al. 2013; Liu et al. 2019). The galaxy was detected in archival radio sky surveys e.g., the Very Large Array Faint Images of the Radio Sky at Twenty-cm (VLA FIRST: Condon et al. 1998). Fig. 1 shows the Pan-STARRS (Chambers et al. 2016) i-band optical image of the galaxy together with the contours of the FIRST radio data (at GHz) and the position of the X-ray counterpart detected by Swift/XRT in . The galaxy is not present in any X-ray catalogs prior to the optical discovery of the transient in .
@@
In addition to the significant IR echo, AT2019aalc shares multi-wavelength characteristics with the two other neutrino candidate transients: a soft X-ray spectrum and transient radio emission (van Velzen et al. 2024). Moreover, similarly to the other two instances, the infrared peak of the transient event is temporarily coincident with the detection time of the likely associated neutrino event. Notably, Winter & Lunardini (2023) estimated the highest neutrino fluence for AT2019aalc among the studied candidates, due to its high estimated SMBH mass (, van Velzen et al. 2024) and low redshift. Since the proposed association with the high-energy neutrino event was announced roughly years after the discovery of the transient event, the source was not at the focus of attention. Thus, the lack of spectroscopic monitoring during the first optical flare hindered a clear classification of the event. Nevertheless, out of the more than AGN detected by the ZTF, less than show similarly rapid and large outbursts (Reusch et al. 2022; van Velzen et al. 2024) suggesting a very unusual transient. Interestingly, roughly years after the first flare, a significant optical re-brightening started in mid-May 2023 (Veres et al. 2023), illustrated by the long-term ZTF difference light curve shown in Fig. 2.
@@ 3.1.1 ZTF
ZTF forced photometry at the location of AT2019aalc was obtained following the recommendations outlined in Masci et al. (2023). We corrected for extinction by adopting a galactic extinction of (Schlafly & Finkbeiner 2011). As shown in Fig. 2, the forced photometry yields a negative flux in the period between the two flares (and also just before the first flare). This implies that during these epochs, the flux decreased below the mean flux of the reference images. For the -band these were obtained between March and May, 2018. While the baseline used to obtain the difference flux of the -band is based on images obtained between March and September, 2018.
@@ 3.1.2 UVOT
The Neil Gehrels Swift Observatory (Arnaud 1996a) performed Target of Opportunity (ToO) observations (project codes: 18972, 19188, 19993, 20997 and 21692, Proposers: Reusch, Veres & Sniegowska) on AT2019aalc from 2023-06-23 to 2025-01-02 with a total exposure time of ks. The Ultraviolet and Optical Telescope (UVOT, Roming et al. 2005) onboard the satellite observed the target field with the following four filters in each epoch: UW2 (central wavelength, 1928 ), UVM2 (2246 ), UW1 (2600 ) and U (3465 ). We analyzed the UVOT images using the Swift UVOT tools included in the heasoft 6.30 software package. We measured the flux using uvotsource, applying a circular aperature of arcseconds (which captures nearly all the flux of the host galaxy). The background was measured using four 6 arcsecond regions in four quadrants away from the target source. To estimate the UV flux before the flare, we use GALEX (Martin et al. 2005) observations obtained in 2007 and 2011. We use the same 15” circular aperture that was applied to the Swift/UVOT observations and use the gPhoton software (Million et al. 2016) to extract the flux. Over the 4 year baseline of the GALEX observations, the NUV magnitudes change by about 0.3 mag. Following van Velzen et al. (2021a), we combine the mean GALEX flux with the SDSS (York et al. 2000) magnitudes and apply Prospector (Johnson & Leja 2017) to find the best-fit stellar population synthesis model (Conroy et al. 2009) that describes these data. From this model we estimate the baseline flux in the UVOT filters and this baseline is subtracted to obtain the difference flux. The baseline flux in the UVM2 band is 17.2 mag, implying an increase of 1.8 mag relative to the peak of the UVM2 light curve. This large increase implies that, at the peak of the flare, we are not sensitive to the value of the baseline flux. This is useful because the baseline of the ZTF forced photometry was obtained around 2018 (at the start of the ZTF survey), while the baseline that is used to obtain the UVOT difference flux is mainly determined by the GALEX observations that were obtained a decade earlier.

### v3

@@ Abstract
Aims. Our aim is to constrain the physical mechanism responsible for the second optical flare, which may also provide clues to the origin of the initial event.
@@
Results. The observations revealed multiple X-ray flares during the second optical flaring episode of the transient and a uniquely bright UV counterpart. The second flare, similarly to the first one, is also accompanied by IR dust echo emission. A long-term radio flare is found with an inverted spectrum. Optical spectroscopic observations reveal the presence of Bowen Fluorescence lines and strong high-ionization coronal lines indicating an extreme level of ionization in the system.
Conclusions. The results suggest that the transient can be classified as a Bowen Fluorescence Flare (BFF), a relatively new sub-class of flaring active galactic nuclei (AGN). AT2019aalc can be also classified as an extreme coronal line emitter (ECLE). We found that, in addition to AT2019aalc, another BFF AT2021loi is spatially coincident with a high-energy neutrino event. We propose a repeating TDE scenario within an AGN framework to explain the multi-wavelength properties of AT2019aalc and suggest a possible connection among ECLEs, BFFs, and TDEs occurring in AGNs.
@@
Tidal disruption events (TDEs) are rare transient events, which occur when a star approaches a supermassive black hole (SMBH) and the tidal forces of the latter rip the star apart (1988Natur.333..523R). About half of the star’s material is accreted around the black hole, generating a luminous outburst across the electromagnetic spectrum. TDE optical light curves typically exhibit a rapid, colorless rise followed by a fast decay on monthly timescales. Several TDEs were detected in X-rays with soft spectra typically explained by late-time accretion disk formation (2020SSRv..216...85S). In addition, a handful of TDEs produced detectable radio emission, explained by delayed111Note that the SMBH accretion rate does not peak on the same timescale as the mass fall-back rate. non-relativistic outflows (delay1; delay2), or only in four instances by newly launched relativistic on-axis jets (2012ApJ...748...36B; 2015MNRAS.452.4297B; 2012ApJ...753...77C; jetted_tdes; 2022cmc). Spectroscopically, TDEs are characterized by a strong blue continuum and broad emission lines ( km s-1) e.g., H and H. In a few cases, Bowen Fluorescence (BF: 1934PASP...46..146B; 1935ApJ....81....1B) lines e.g., N iii at appear (e.g., 2019ApJ...873...92B; spectra2).
@@
TDEs generally produce one optical flare, however, in a few cases a re-brightening episode was detected, even years after the initial flare. These might be explained in the following way. A star on a grazing orbit might be partially disrupted and only a fraction of the stellar mass will be deposited onto the SMBH. The secondary disruption of a surviving core after it reaches the pericenter again results in re-brightening. This scenario is naturally explained by the Hills capture mechanism (1988Natur.331..687H), where a binary star system is disrupted by the SMBH, capturing one star on a highly eccentric orbit prone to repeated partial disruptions. In this process, the captured star’s orbit brings it close enough to the SMBH multiple times, causing episodic stripping of stellar material and producing recurring flares over timescales of months to years. Only a few partial TDEs or candidates are known to date. These are AT2020vdq (2020vdq), ASASSN14ko (2021ApJ...910..125P; 14ko2; 14ko3; 2023ApJ...956L..46H), eRASSt J045650.3203750 (2023A&A...669A..75L; 2024A&A...683L..13L), RX J133157.6-324319.7 (2022RAA....22e5004H), AT2018fyk (fyk1; fyk2; 2024ApJ...970..116W; 2024ApJ...971L..31P), AT2019avd (avd1; 2021ApJ...920...56F; avd2; avd3; avd4), AT2019azh (2021MNRAS.500.1673H), AT2022dbl (2024ApJ...971L..26L; 2024arXiv241215326H; 2025ApJ...987L..20M) and AT2021aeuk (at2021aeuk_jingbo).
@@
Based on the observed properties of TDE host galaxies, 2014ApJ...792...53V estimated a TDE rate of a few per year per galaxy, which is well borne out by recent data (2023ApJ...955L...6Y) and earlier observational estimates (1999MNRAS.306...35S). Importantly, a similar fraction of TDEs should occur in AGN hosts (e.g., tdes_in_agn; 2024MNRAS.527.8103R). These transients are challenging to distinguish from regular AGN variability and only a few candidate TDEs occurring in AGN (in the following: TDE-AGN) have been studied earlier (e.g., PS16dtm - a TDE in a narrow-line Seyfert-1 (NLSy1) galaxy (2017ApJ...843..106B) or the multiple soft X-ray flares in IC 3599 (2015A&A...581A..17C) and GSN 069 (2018ApJ...857L..16S)). In contrast with TDEs seen in earlier quiescent galaxies, these transients are characterized by slowly decaying optical emission and bumps tend to appear after the peak. It is worth noting that transients classified as Ambiguous Nuclear Transients (ANTs; e.g., 2025MNRAS.537.2024W), as well as several of the events studied by 2021ApJ...920...56F, could also be TDEs in AGN hosts, potentially increasing their overall number.
@@
Recently, Trakhtenbrot identified a new class of spectroscopically unique flares from accreting SMBHs. In , the transient events AT2017bgt, F01004-2237 and OGLE17aaj were identified as the first members of the class based on their unique optical spectroscopic features. Later, the peculiar X-ray transient AT2019avd (avd1; 2021ApJ...920...56F; avd2; avd3; avd4) was found with remarkable common properties, however, its classification as an AT2017bgt-like transient is unclear. A more promising case is AT2021loi (AT2021loi), a UV-bright transient event with very similar multi-wavelength properties to the three flares studied by Trakhtenbrot. These events show unobscured AGN-like optical spectra with significant and persistent BF lines, high UV-to-optical ratio and optical emission that decays on yearly timescales. These AGN are characterized by accretion that intensifies significantly over long timescales. Hereafter, we will refer to this new class of SMBH flares as Bowen Fluorescence Flares (BFFs), following AT2021loi. Using the Zwicky Transient Facility (ZTF, 2019PASP..131a8003M) Public Survey data, 2023ApJ...957...57D found only one BFF candidate (AT2021seu) among spectroscopically classified transients, suggesting that these flares are rare. Two other candidates, namely AT2022fpx, (2024MNRAS.532..112K) and AT2020afhd (2024TNSAN..53....1A) were found later.
@@
In recent years, two nuclear transients were found to be spatially coincident with neutrino events detected by the IceCube Neutrino Observatory based on ZTF follow-up campaigns of IceCube neutrino alerts (2023MNRAS.521.5046S). The TDE AT2019dsg is likely associated with the neutrino, IC191001A (Bran) and a further promising association, between the candidate TDE AT2019fdr and IC200530A (Tywin) points to TDEs as a potential new class of sources of the extragalactic neutrinos. Both sources were identified as having luminous mid-infrared emission, and a search for similar flares (Lancel) led to the identification of the transient event AT2019aalc (a.k.a. ZTF19aaejtoy) with the neutrino alert IC191119A. In addition, two candidate obscured TDEs (hidden) and the extremely energetic candidate TDE AT2021lwx (lwx) were associated with high-energy neutrinos. A recently claimed association between a neutrino flare detected by the IceCube Observatory and an X-ray flare from the TDE ATLAS17jrp suggests that the X-ray emission of these sources plays a crucial rule in producing neutrinos (xray_neu).
@@
In this paper, we focus on the nuclear transient event and neutrino source candidate AT2019aalc, especially on the significant optical re-brightening that started four years after its initial flare (re-br).
@@
The paper is organized as follows. We discuss the discovery of the transient event and give an insight about its host galaxy in Sect. 2, summarize our observing campaign and the data reduction in Sect. 3, present the results of the observations in Sect. 4, discuss them in Sect. 5 and finally give a summary and conclude in Sect. 6. Throughout the paper, we adopt a flat CDM cosmological model with parameters , , and . In this model (2006PASP..118.1711W), mas angular size corresponds to pc projected linear size at the source redshift (2012ApJS..203...21A). This redshift corresponds to a luminosity distance of Mpc. All magnitudes are given in AB system (1974ApJS...27...21O) unless stated otherwise.
@@ 2 The transient event: AT2019aalc
An analysis of the infrared properties of AT2019dsg and AT2019fdr revealed a strong reverberation signal, particularly an the infrared dust echo, stemming from reprocessed emission of optical/UV photons to the infrared by interaction with hot dust around the SMBH, at a distance of pc (2016ApJ...828L..14J; 2016ApJ...832..188D; 2016MNRAS.458..575L; 2021SSRv..217...63V; 2021ApJ...911...31J; Lancel). In addition to the exceptionally high infrared luminosities of these transients, the neutrino detection times appear to be temporally consistent with the infrared luminosity peaks (Bran; Tywin). This discovery motivated a systematic search for neutrino emission using an extended sample of black hole flares (Lancel). This archival search revealed the nuclear transient AT2019aalc, with a powerful optical flare and dust echo emission in spatial coincidence with the high-energy neutrino event, IC191119A (2019GCN.26258....1I), detected by the IceCube Observatory.
@@
The host galaxy of AT2019aalc (2MASX J15241665+0451192) is a barred spiral (SBa) galaxy hosting an active nucleus (2013AJ....146..151O). Based on its Sloan Digital Sky Survey (SDSS: sdssdr7) spectrum it was further classified as a broad-line Seyfert-1 galaxy (2013ApJ...763..133T; 2019ApJS..243...21L). The galaxy was detected in archival radio sky surveys e.g., the Very Large Array Faint Images of the Radio Sky at Twenty-cm (VLA FIRST: 1998AJ....115.1693C). Fig. 1 shows the Pan-STARRS (2016arXiv161205560C) -band optical image of the galaxy, together with the contours of the FIRST radio data (at GHz) and the position of the X-ray counterpart detected by the Swift/XRT in . The galaxy is not present in any X-ray catalogs prior to the optical discovery of the transient in . In addition to the significant IR echo, AT2019aalc shares multi-wavelength characteristics with the two other neutrino candidate transients: a soft X-ray spectrum and transient radio emission (Lancel). Moreover, similarly to the other two instances, the infrared peak of the transient event is temporarily coincident with the detection time of the likely associated neutrino event. Notably, 2023ApJ...948...42W estimated the highest neutrino fluence for AT2019aalc among the studied candidates, due to its high estimated SMBH mass (, Lancel) and low redshift. Since the proposed association with the high-energy neutrino event was announced roughly two years after the discovery of the transient event, the source was not at the focus of attention. Thus, the lack of spectroscopic monitoring during the first optical flare hindered a clear classification of the event. Nevertheless, out of the more than AGN detected by the ZTF, less than show similarly rapid and large outbursts (Tywin; Lancel), suggesting a very unusual transient. Interestingly, roughly four years after the first flare, a significant optical re-brightening started in mid-May 2023 (re-br), illustrated by the long-term ZTF difference light curve shown in Fig. 2.
@@ 3.1.1 ZTF
ZTF forced photometry at the location of AT2019aalc was obtained following the recommendations outlined in 2019PASP..131a8003M. We corrected for extinction by adopting a galactic extinction of (2011ApJ...737..103S). As shown in Fig. 2, the forced photometry yields a negative flux in the period between the two flares (and also just before the first flare). This implies that during these epochs, the flux decreased below the mean flux of the reference images. For the -band, these were obtained between March 2018 and May 2018. While the baseline used to obtain the difference flux of the -band is based on images obtained between March and September, 2018.
@@ 3.1.2 UVOT
The Neil Gehrels Swift Observatory (2004ApJ...611.1005G) performed Target of Opportunity (ToO) observations (project codes: 18972, 19188, 19993, 20997 and 21692, Proposers: Reusch, Veres & Sniegowska) on AT2019aalc from 2023-06-23 to 2025-01-02 with a total exposure time of ks. The Ultraviolet and Optical Telescope (UVOT, uvot) onboard the satellite observed the target field with the following four filters in each epoch: UVW2 (central wavelength, 1928 ), UVM2 (2246 ), UVW1 (2600 ) and U (3465 ). We analyzed the UVOT images using the Swift/UVOT tools included in the heasoft 6.30 software package. We measured the flux using uvotsource, applying a circular aperture of arcseconds (which captures nearly all the flux of the host galaxy). The background was measured using four 6 arcsecond regions in four quadrants away from the target source. To estimate the UV flux before the flare, we use GALEX (Martin05) observations obtained in 2007 and 2011. We use the same 15” circular aperture that was applied to the Swift/UVOT observations and use the gPhoton software (Million16) to extract the flux. Over the four year baseline of the GALEX observations, the NUV magnitudes change by about 0.3 mag. Following vanVelzen20, we combine the mean GALEX flux with the SDSS (york02) magnitudes and apply Prospector (Johnson17) to find the best-fit stellar population synthesis model (Conroy09) that describes these data. From this model we estimate the baseline flux in the UVOT filters and this baseline is subtracted to obtain the difference flux. The baseline flux in the UVM2 band is 17.2 mag, implying an increase of 1.8 mag relative to the peak of the UVM2 light curve. This large increase implies that, at the peak of the flare, we are not sensitive to the value of the baseline flux. This is useful because the baseline of the ZTF forced photometry was obtained around 2018 (at the start of the ZTF survey), while the baseline that is used to obtain the UVOT difference flux is mainly determined by the GALEX observations that were obtained a decade earlier.